Title

INAF queue observing, 2014 Oct 29-30

Observer: E. Sani, R. Carini
Telescope Operator: S. Allanson
Support Astronomer: M. Ewards on call
Instrument: LUCI, MODS, LBC-PARALLEL

Summary

When the Dome is opened the humidity is 8%, no wind, the sky is clear, seeing 0.6

Data Summary

Project_RUN Objects Filter ExpTime Notes
Cresci DSF2237b Ks 2h PROGRAM COMPLETED
Standard feige110 dual and red grating   airmass 1.5
Saracco F XLLSM1 red grating + LBC red 2 h 30 m COMPLETED
Dotto KC46 V- ,R -Bessel 1 h 15m PROGRAM COMPLETED
Clementini IZw 18 V- I- Bessel 10m  

Details

01:17 Flat for LUCI imaging luci_20141030_0001-0016
  • luci_20141030_0005-0016 are good

01:32 slew to scientific target of CRESCI program

01:41 run executeLUCIScript.sh CampoB _Ks.img luci_20141030_0017
  • the guide star is too faint

01:49 run executeLUCIScript.sh CampoB _Ks.img with another guide star luci_20141030_0018-0136
  • average seeing 0.6

05:23 slew to the standard star FS 152 executeLUCIScript.sh F2-152_Ks.img
  • the camera in the guider stops working, the LUCI machine wasn't able to comunicate with guider. The TO restarted it.

05:30 executeLUCIScript.sh F2-152_Ks.img
  • same problem with the LUCI machine

05:34 executeLUCIScript.sh F2-152_Ks.img luci_20141030_0137-0141
  • since the seeing is very good (0.5), the collimation is not perfect, in the first images the stars are distorted.
  • we re-run the script with two offsets luci_20141030_0142-0143

05.42 we ask to TO to reconfigure the parallel mode.

06:00 LBC problem: the red channel: one or more systems failed.
  • filters have problems on power line
  • The TO call the SA

06.07 run the acquisition of the standard with MODS
  • acqMODS feige110_rec.acq (recovery procedure) mods1r_20141030_ 0001-0002
  • through slit image is blurry due because the sistem was still collimating, anthow the star is well within the slit.

06:19 execMODS feige110.obs mods1r.0003-0008, mods1b.0001-0003

06:50 acqMODS mods.5.xllsm1_recover.acq mods1r.0009-00010

06:56 dofpia, /X2, /first, /redonly on the scientific field Saracco_F

07.00 execMODS mods.5.xllsm1_6.obs mods1r.00011-014 and run XLSSm1_V_180s.ob
  • lbcr.20141030.070046 is blurred
  • the mean FWHM on the GCS is 0.74, and 0.75 measure on LBCred images
  • lbcr.071092 is blurred

07.35 dofpia, /X2, /redonly on field XLSSm1_I_180s.ob

07.41 Run XLSSm1_V_180s.ob * lbcr.075042 is blurred cause by offset of MODS

08.09 execMODS mods.5.xllsm1_7.obs =mods1r.00015-018

08.20 dofpia, /X2, /redonly

08.25 run XLSSm1_V_180s.ob
  • lbcr.082941 is blurred caused by offset of MODS
  • lbcr.084427 is blurres caused by offset of MODS

09.05 dofpia, /X2, /redonly

09.08 run XLSSm1_V_180s.ob

09:20 we ask to TO to switch to the bimodal LBC * TO clears both primary optics

09.42 we pass to Dotto program
  • dofpia, /X2 /first on copointing star WT10-271 lbcr.095518,lbcb.095535

10.00 problem with offsets too big
  • we try again

10:04 slew to the scientific target, run 2014
  • copointing failed. To don't waist time we try to recover the wrong copointing offsets running lbcrangebal on the images already acquires. This fails because the telescope lost the preset.
  • we have tried again the copointing but failed we have to do the dofpia because the TO have to clear all the commands committed to telescope and active optics as well as the pointing settings

10.10 dofpia, /X2 /first on copointing star WT10-271

10:21 run the copointing star WT10-271 lbcr.102202,lbcb.102205

10:25 run 2014KC46.ob
  • we lost one image in the blue channel because it had no filter. We have to kill the preset and acquired again the ob.

12:03 slew to Clementini's program IZw18

12.06 run dofpia , /X2

12:10 running copointing

12:16 run IZ18_4.ob

12:30 We star skyflats
  • Run B- V-Bessel

13:20 The TO closes the Dome
  • we run LBC Bias
  • we run mods long slit (5 arcsec) flats dual grating with lamp QTH + ND1.5 modsb1.0004-0006

-- %USERSIG{RobertaCarini - 2014-10-30}%

Comments

 
Topic revision: r7 - 02 Nov 2014, EleonoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback